# DeepAgent: A General Reasoning Agent with Scalable Toolsets

### v1

@@ Abstract.
journalyear: 2026doi: xxxxxxx.xxxxxxx
1. Introduction
The rapid advancement of large language models (LLMs) has inspired the development of LLM-powered agents, which have found broad applications in scenarios such as web information seeking, software engineering, and personal assistance (Wang et al., 2024b; Qu et al., 2025b; Jin et al., 2024). Existing agent frameworks predominantly rely on predefined workflows, exemplified by methods like ReAct (Yao et al., 2022b) and Plan-and-Solve (Wang et al., 2023), which employ structured planning processes and iterative “Reason-Act-Observe” cycles as illustrated in Figure 2(a). Although effective in simpler tasks, these approaches suffer from several critical limitations: (1) lack of autonomy in execution steps and overall procedure; (2) inability to dynamically discover tools during task execution; (3) deficiency in fully autonomous management of interactive memory; and (4) insufficient depth and coherence in reasoning about the entire task. These fundamental shortcomings severely constrain the agents’ ability to tackle real-world problems, particularly for complex tasks that demand general tool-use and long-horizon interaction with the environment.
Recently, the advent of large reasoning models (LRMs) has demonstrated the capability to solve complex problems in domains like mathematics, programming, and scientific reasoning through a step-by-step “slow thinking” process (Chen et al., 2025b; Li et al., 2025e). However, many real-world tasks necessitate the use of external tools for their completion. While some studies have explored new paradigms for integrating tool use within the reasoning process, such as Search-o1 (Li et al., 2025a), DeepResearcher (Zheng et al., 2025a), and ToRL (Li et al., 2025f), these approaches are often restricted to a limited set of predefined tools, such as web search, page browsing, and coding (Figure 2(b)). This constrained set of tools significantly hinders their applicability to a wide range of complex, real-world scenarios.
@@ 1. Introduction
To address these challenges, we introduce DeepAgent, an end-to-end deep reasoning agent that can complete an entire task by dynamically retrieving and calling tools within a single, coherent agentic reasoning process. As depicted in Figure 2(c), DeepAgent operates by autonomously thinking, searching for tools, and executing actions. This paradigm shifts away from traditional, predefined workflows that rely on predefined tools, task planning, and iterative tool use, where each generation step focuses only on the immediate objective. Instead, DeepAgent maintains a global perspective on the entire task, unconstrained by the need to deliberate on specific, isolated operations. Tools are not pre-retrieved in advance but are dynamically discovered on an as-needed basis, thereby fully unlocking the autonomous potential of the large reasoning model.
To empower DeepAgent to thoroughly and robustly explore new tools and navigate complex environments during long-horizon interactions, we equip it with memory management capabilities. We introduce an Autonomous Memory Folding strategy that allows DeepAgent to consolidate its previous thoughts and interaction history into a structured memory schema at any point during its thinking before resuming the agentic reasoning process. This mechanism not only saves tokens and enhances reasoning efficiency over extended interactions but also provides the agent an opportunity to “take a breath”, preventing it from becoming trapped in wrong exploration paths and enabling it to reconsider its strategy, thus improving the overall success rate. To mitigate information loss during this folding process, we introduce a brain-inspired memory architecture comprising episodic memory, working memory, and tool memory, all structured with an agent-usable data schema to ensure the stability and utility of the compressed memory.
@@ 2.1. Large Reasoning Models
Large Reasoning Models (LRMs) (DeepSeek-AI et al., 2025; Jaech et al., 2024) have demonstrated significant performance improvements in mathematical, scientific, and coding tasks by employing step-by-step slow thinking processes before generating final responses. Existing research has explored various approaches to elicit extended Chain-of-Thought (CoT) reasoning (Wei et al., 2022) from models, including data synthesis for Supervised Fine-Tuning (SFT) (Qin et al., 2024a; Min et al., 2024; Ye et al., 2025b), and end-to-end RL (DeepSeek-AI et al., 2025; Hu et al., 2025). Additionally, substantial work has investigated optimization strategies for reasoning models, such as advanced RL training algorithms (Yu et al., 2025; Zheng et al., 2025b) and improving reasoning efficiency (Chen et al., 2024; Yang et al., 2025b). However, models relying solely on parametric knowledge face inherent limitations and cannot interact with the real world. Recent studies have begun exploring tool-augmented reasoning approaches, including Search-o1 (Li et al., 2025a), Search-R1 (Jin et al., 2025b), ToRL (Li et al., 2025f), DeepResearcher (Zheng et al., 2025a), and SimpleTIR (Xue et al., 2025). However, these methods typically support only a limited set of research-oriented tools, such as web search, page browsing, and code execution, which constrains their applicability to real-world scenarios that demand access to more diverse tools.
@@ 2.2. Autonomous Agents
LLM-powered autonomous agents accomplish real-world tasks by invoking external tools to interact with their environment (Wang et al., 2024b; Schick et al., 2023; Qu et al., 2025a; Zhang et al., 2025; Wu et al., 2025; Li et al., 2025d; Dong et al., 2025b; Huang et al., 2025; Wang et al., 2025c; Liu et al., 2025a; Jin et al., 2025c). Current agent methodologies, including ReAct (Yao et al., 2022b), Plan-and-Solve (Wang et al., 2023), Reflextion (Shinn et al., 2024), and CodeAct (Wang et al., 2024a), predominantly follow predefined workflows with fixed execution patterns. This rigid structure limits their ability to fully leverage the autonomous decision-making and deep reasoning capabilities of advanced reasoning models. Recent efforts have investigated training LLMs to autonomously invoke tools through data synthesis and SFT methods (Su et al., 2025; Fang et al., 2025; Xiao et al., 2025) and RL training frameworks (Feng et al., 2025; Wang et al., 2025b; Jiang et al., 2025; Xi et al., 2025; Liu et al., 2025b; Gao et al., 2025; Kang et al., 2025; Sun et al., 2025; Dong et al., 2025c; Chen et al., 2025a; Dong et al., 2025a; Li et al., 2025c). However, most existing methods rely on pre-selected, labeled tools, which limit their applicability to real-world scenarios. Real-world tasks are highly variable and require access to diverse toolsets that cannot be predetermined, aligning with the emerging Model Context Protocol (MCP) (Hou et al., 2025) paradigm. Although some prior work has explored tool retrieval mechanisms (Qin et al., 2024b; Wang et al., 2025a; Shi et al., 2025), most approaches conduct only a single upfront retrieval step and incorporate the retrieved tools, with limited exploration of dynamic tool discovery during task execution. Therefore, we aim to develop a deep reasoning agent capable of dynamically discovering and invoking helpful tools from scalable toolsets to address more generalized real-world tasks.
@@ 3. Methodology
In this section, we first formulate the task of autonomous agentic reasoning. Then, we provide a detailed overview of the DeepAgent framework. Finally, we elaborate on the core components of DeepAgent, including the mechanism for autonomous tool use and memory folding, the brain-inspired memory schema, and our end-to-end reinforcement learning training method, ToolPO.
@@ 3.1. Problem Formulation
  • Tool Search ($a_{t}^{\text{search}}$): A natural language query $q_{s}$ to find relevant tools from the toolset $\mathcal{T}$. The observation $o_{t}$ is a list of retrieved tools.
@@
The sequence of states, actions, and observations forms a trajectory $\tau=(s_{1},a_{1},o_{1},\dots,s_{T},a_{T},o_{T})$. The process terminates when the agent completes the task or reaches a maximum step limit. The objective is to learn an optimal policy $\pi_{\theta}^{*}$ that maximizes the expected cumulative reward for a given task:
(2) $\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$
where $R(\tau)$ is a reward function that evaluates the overall success of the trajectory $\tau$.
@@ 3.3. Autonomous Tool Search and Calling
(3) $\mathcal{T}_{\text{retrieved}}=\underset{\tau_{i}\in\mathcal{T}}{\text{top-k}}\left(\text{sim}(E(q_{s}),E(d_{i}))\right).$
@@ 3.4. Autonomous Memory Folding and Brain-Inspired Memory Schema
To ensure that the compressed memory is stable and easily parsed by the agent, we employ an agent-usable data schema in JSON format instead of unstructured natural language. This structured format offers two main benefits: it maintains a controllable and predictable structure, and it mitigates the loss of critical details that can occur when summarizing long-form text. Details of the data schema are provided in Appendix D.
@@ 3.5. End-to-end RL Training with ToolPO
For each input prompt, we sample a group of $K$ trajectories $\{\tau_{1},\dots,\tau_{K}\}$. ToolPO defines two distinct reward components. The first is a reward for overall task success, $R_{\text{succ}}(\tau)$, which is a task success score reflecting the quality of the final outcome (e.g., the accuracy of the final answer). The second is a tool-call reward, $R_{\text{action}}(\tau)$, which reflects the quality of intermediate actions. This action-level reward is composed of rewards for correct tool invocations and efficient memory folding. Specifically, $R_{\text{action}}(\tau)=\lambda_{1}\sum_{t=1}^{T}C(a_{t}^{\text{call}})+\lambda_{2}S_{\text{pref}}(\tau)$, where $C(a_{t}^{\text{call}})$ is 1 if a tool call is correct and 0 otherwise. $S_{\text{pref}}(\tau)$ is a preference score encouraging efficient use of memory folding, defined by comparing a trajectory with folding ($\tau_{\text{fold}}$) to one without ($\tau_{\text{direct}}$): $S_{\text{pref}}=(L(\tau_{\text{direct}})-L(\tau_{\text{fold}}))/(L(\tau_{\text{direct}})+L(\tau_{\text{fold}}))$.
@@
(5) $A_{\text{succ}}(\tau_{k})=R_{\text{succ}}(\tau_{k})-\frac{1}{K}\sum\nolimits_{j=1}^{K}R_{\text{succ}}(\tau_{j}).$
This advantage is attributed to all generated tokens in the trajectory, providing a global learning signal. Similarly, the action-level advantage is:
@@
$\mathbb{E}_{\tau_{k}}\left[\sum\nolimits_{i=1}^{|\tau_{k}|}\min\Big(\rho_{i}(\theta)A(y_{i}),\text{clip}(\rho_{i}(\theta),1-\epsilon,1+\epsilon)A(y_{i})\Big)\right],$
@@ 4.1. Tasks and Datasets
General Tool-Use. These benchmarks encompass a broad range of distinct tools, scaling from tens to over ten thousand, making them ideal for evaluating the scalability of different approaches. We utilize four representative scenarios: ToolBench (Qin et al., 2024b), based on over 16,000 real-world APIs, for which we use the G3 subset requiring multi-step, multi-tool calls; API-Bank (Li et al., 2023), which includes 314 human-annotated dialogues with 73 APIs and 753 API calls, to assess planning, retrieval, and calling capabilities; RestBench (Song et al., 2023), comprising scenarios from the TMDB movie database (54 tools, avg. 2.3 calls/question) and the Spotify music player (40 tools, avg. 2.6 calls/question) to simulate typical REST applications; and ToolHop (Ye et al., 2025a), a multi-hop reasoning dataset with 3,912 locally executable tools that necessitate 3 to 7 sequential tool calls per task. For these tasks, we adopt two settings: given ground-truth tools and given entire toolsets with tool retrieval capabilities.
Downstream Applications. We evaluate our approach on several downstream applications that require domain-specific toolsets. These include ALFWorld (Shridhar et al., 2021), a text-based embodied AI task where agents complete goals using nine basic actions (e.g., move, take); WebShop (Yao et al., 2022a), an online shopping environment with ‘search’ and ‘click’ actions to fulfill users’ specific product purchasing requirements; GAIA (Mialon et al., 2024), a complex information-seeking benchmark where we equip the agent with tools for web search, page browsing, Visual Question Answering (VQA), code compilation, and file reading; and Humanity’s Last Exam (HLE) (Phan et al., 2025), a set of highly difficult reasoning problems, for which we provide code, search, page browsing, and VQA tools. These benchmarks test the agent’s ability to perform long-horizon planning and robust interaction in complex, real-world scenarios. For this category of tasks, we provide agents with task-specific toolsets.
@@ 4.2. Baselines
Our baselines include: (1) Workflow-based Methods: ReAct (Yao et al., 2022b) alternates explicit reasoning with environment actions in a Reason-Act-Observe loop. CodeAct (Wang et al., 2024a) expresses actions as executable Python code that runs in an interpreter. Plan-and-Solve (Wang et al., 2023) first sketches a high-level plan and then executes it step by step. Reflexion (Shinn et al., 2023) enhances learning through verbal self-reflection after failed attempts. AgentLM (Zeng et al., 2024) uses instruction tuning to enhance general agent capabilities of LLMs. (2) Autonomous Tool Usage within Reasoning: WebThinker (Li et al., 2025b) interleaves thinking with web search and deep web exploration. HiRA (Jin et al., 2025a) introduces a hierarchical agent architecture where a meta planner decomposes tasks, a coordinator routes subtasks, and specialized executors solve them with dual-channel memory. OpenAI Deep Research (OpenAI, 2025) is an agentic system based on reasoning models.

### v2

@@ Abstract.
$\dagger$ Corresponding author.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792460ccs: Computing methodologies Planning and scheduling
@@
1. Introduction
The rapid advancement of Large Language Models (LLMs) has inspired the development of LLM-powered agents, which have found broad applications in scenarios such as web information seeking, software engineering, and personal assistance (Qu et al., 2025b; Jin et al., 2024). Most existing agents follow predefined workflows (e.g., ReAct (Yao et al., 2022b) and Plan-and-Solve (Wang et al., 2023)) with iterative “Reason-Act-Observe” loops (Figure 2(a)). Although effective in simpler tasks, these approaches suffer from several critical limitations: (1) lack of autonomy in execution steps and overall procedure; (2) inability to dynamically discover tools during task execution; (3) deficiency in fully autonomous management of interactive memory; and (4) insufficient depth and coherence in reasoning about the entire task. These limitations hinder agents from real-world problems, particularly for complex tasks that demand general and multiple tool-use.
Recently, the advent of Large Reasoning Models (LRMs) has demonstrated the capability to solve complex problems in domains like mathematics, programming, and scientific reasoning through a step-by-step “slow thinking” process (Chen et al., 2025a; Li et al., 2025e; Wang et al., 2025c). However, many real-world tasks necessitate the use of external tools for their completion. Recent approaches (e.g., Search-o1 (Li et al., 2025a), DeepResearcher (Zheng et al., 2025), and ToRL (Li et al., 2025d)) integrate tool use into reasoning, but typically rely on a small, fixed tool set such as search, browsing, and coding (Figure 2(b)), limiting their generality.
To address these challenges, we introduce DeepAgent, an end-to-end deep reasoning agent that can complete an entire task by dynamically retrieving and calling tools within a single, coherent agentic reasoning process. As depicted in Figure 2(c), DeepAgent operates by autonomously thinking, searching for tools, and executing actions. This paradigm shifts away from traditional, predefined workflows that rely on predefined tools, task planning, and iterative tool use. Instead, DeepAgent maintains a global perspective on the entire task, unconstrained by the need to deliberate on specific, isolated operations. Tools are not pre-retrieved in advance but are dynamically discovered on an as-needed basis, thereby fully unlocking the autonomous potential of the large reasoning model.
To empower DeepAgent to thoroughly and robustly explore new tools and navigate complex environments during long-horizon interactions, we equip it with memory management capabilities. We introduce an Autonomous Memory Folding strategy that allows DeepAgent to consolidate its previous thoughts and interaction history into a structured memory schema at any point during its thinking before resuming the agentic reasoning process. This mechanism not only saves tokens and enhances reasoning efficiency over extended interactions but also provides the agent an opportunity to “take a breath”, preventing it from becoming trapped in wrong exploration paths and enabling it to reconsider its strategy. To mitigate information loss during folding, we introduce a brain-inspired memory architecture comprising episodic memory, working memory, and tool memory, all structured with an agent-usable data schema to ensure the stability and utility of the folded memory.
@@ 2.1. Large Reasoning Models
Large Reasoning Models (LRMs) (DeepSeek-AI et al., 2025; Jaech et al., 2024) have demonstrated significant performance improvements in mathematical, scientific, and coding tasks by employing step-by-step slow thinking processes before generating final responses. Existing research has explored various approaches to elicit extended Chain-of-Thought (CoT) reasoning (Wei et al., 2022) from models, including data synthesis for Supervised Fine-Tuning (SFT) (Qin et al., 2024a; Wang et al., 2025b), and end-to-end RL (DeepSeek-AI et al., 2025). Additionally, substantial work has investigated optimization strategies for reasoning models, such as advanced RL training algorithms (Wang et al., 2024b) and improving reasoning efficiency (Yang et al., 2025b). However, models relying solely on parametric knowledge face inherent limitations and cannot interact with the real world. Recent studies have begun exploring tool-augmented reasoning approaches, including Search-o1 (Li et al., 2025a), Search-R1 (Jin et al., 2025a), ToRL (Li et al., 2025d), DeepResearcher (Zheng et al., 2025), and SimpleTIR (Xue et al., 2025). However, these methods typically support only a limited set of research-oriented tools, such as web search, page browsing, and code execution, which constrains their applicability to real-world scenarios that demand access to more diverse tools.
@@ 2.2. Autonomous Agents
LLM-powered autonomous agents accomplish real-world tasks by invoking external tools to interact with their environment (Qu et al., 2025a; Wu et al., 2025; Dong et al., 2025b; Liu et al., 2025a; Li et al., 2024; Jin et al., 2025d; Li et al., 2025c; Dong et al., 2025c; Jin et al., 2025b; Tan et al., 2025; Wang et al., 2026; Zhang et al., 2025; Song et al., 2026; Hu et al., 2024, 2025b, 2025a). Current agent methodologies, including ReAct (Yao et al., 2022b), Plan-and-Solve (Wang et al., 2023), Reflextion (Shinn et al., 2024), and CodeAct (Wang et al., 2024a), predominantly follow predefined workflows with fixed execution patterns. This rigid structure limits their ability to fully leverage the autonomous decision-making and deep reasoning capabilities of advanced reasoning models. Recent efforts have investigated training LLMs to autonomously invoke tools through data synthesis and SFT methods (Fang et al., 2025; Xiao et al., 2025) and RL training frameworks (Feng et al., 2025; Wang et al., 2025d; Jiang et al., 2025; Liu et al., 2025b; Kang et al., 2025; Sun et al., 2025; Dong et al., 2025d; Chen et al., 2025b; Dong et al., 2025a). However, most existing methods rely on pre-selected, labeled tools, which limit their applicability to real-world scenarios. Real-world tasks are highly variable and require access to diverse toolsets that cannot be predetermined, aligning with the emerging Model Context Protocol (MCP) (Hou et al., 2025) paradigm. Although some prior work has explored tool retrieval mechanisms (Qin et al., 2024b; Wang et al., 2025a; Shi et al., 2025), most approaches conduct only a single upfront retrieval step and incorporate the retrieved tools, with limited exploration of dynamic tool discovery during task execution. Therefore, we aim to develop a deep reasoning agent capable of dynamically discovering and invoking helpful tools from scalable toolsets to address more generalized real-world tasks.
@@ 3.1. Problem Formulation
  • Tool Search ($a_{t}^{\text{search}}$): A natural language query $q_{s}$ to find relevant tools from $\mathcal{T}$. The observation $o_{t}$ is a list of retrieved tools.
@@
The sequence of states, actions, and observations forms a trajectory $\tau=(s_{1},a_{1},o_{1},\dots,s_{T},a_{T},o_{T})$. The process terminates when the agent completes the task or reaches a maximum step limit. Suppose $R(\tau)$ is a reward function that evaluates the overall success of the trajectory $\tau$, the objective is to learn an optimal policy $\pi_{\theta}^{*}$ that maximizes the expected cumulative reward for a given task:
(2) $\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)].$
@@ 3.3. Autonomous Tool Search and Calling
(3) $\mathcal{T}_{\text{retrieved}}=\underset{\tau_{i}\in\mathcal{T}}{\text{top-k}}\left(\text{sim}\left(E\left(q_{s}\right),E\left(d_{i}\right)\right)\right).$
@@ 3.4. Autonomous Memory Folding and Brain-Inspired Memory Schema
To ensure that the compressed memory is stable and easily parsed by the agent, we employ an agent-usable data schema in JSON format instead of unstructured natural language. It offers two main benefits: maintaining a controllable and predictable structure, and mitigating the loss of critical details that can occur when summarizing long-form text. Details of the data schema are in Appendix C.
@@ 3.5. End-to-end RL Training with ToolPO
For each input prompt, we sample a group of $K$ trajectories $\{\tau_{1},\dots,\tau_{K}\}$. ToolPO defines two distinct reward components. The first is a reward for overall task success, $R_{\text{succ}}(\tau)$, which is a task-success score that reflects the quality of the final outcome (e.g., the accuracy of the final answer). The second is a tool-call reward, $R_{\text{action}}(\tau)$, which reflects the quality of intermediate actions. This action-level reward is composed of rewards for correct tool invocations and efficient memory folding. Specifically, $R_{\text{action}}(\tau)=\lambda_{1}\sum_{t=1}^{T}C(a_{t}^{\text{call}})+\lambda_{2}S_{\text{pref}}(\tau)$, where $C(a_{t}^{\text{call}})$ is 1 if a tool call is correct and 0 otherwise. $S_{\text{pref}}(\tau)$ is a preference score encouraging efficient use of memory folding, defined by comparing a trajectory with folding ($\tau_{\text{fold}}$) to one without ($\tau_{\text{direct}}$): $S_{\text{pref}}=(L(\tau_{\text{direct}})-L(\tau_{\text{fold}}))/(L(\tau_{\text{direct}})+L(\tau_{\text{fold}}))$.
@@
(5) $A_{\text{succ}}(\tau_{k})=R_{\text{succ}}(\tau_{k})-\frac{1}{K}\sum\nolimits_{j=1}^{K}R_{\text{succ}}(\tau_{j}),$
which is attributed to all generated tokens in $\tau_{k}$, providing a global learning signal. Similarly, the action-level advantage is:
@@
$\mathbb{E}_{\tau_{k}}\left[\sum\nolimits_{i=1}^{|\tau_{k}|}\min\Big(\rho_{i}(\theta)A(y_{i}),\text{clip}(\rho_{i}(\theta),1-\epsilon,1+\epsilon)A(y_{i})\Big)\right].$
@@ 4.1. Tasks and Datasets
General Tool-Use. These benchmarks encompass a broad range of distinct tools, scaling from tens to over ten thousand, making them ideal for evaluating the scalability of different approaches. We utilize four representative scenarios: ToolBench (Qin et al., 2024b), based on over 16,000 real-world APIs, for which we use the G3 subset requiring multi-step, multi-tool calls; API-Bank (Li et al., 2023), which includes 314 human-annotated dialogues with 73 APIs and 753 API calls, to assess planning, retrieval, and calling capabilities; RestBench (Song et al., 2023), comprising scenarios from the TMDB movie database (54 tools, avg. 2.3 calls/question) and the Spotify music player (40 tools, avg. 2.6 calls/question) to simulate typical REST applications; and ToolHop (Ye et al., 2025), a multi-hop reasoning dataset with 3,912 locally executable tools that necessitate 3 to 7 sequential tool calls per task. For these tasks, we adopt two settings: given ground-truth tools and given entire toolsets with tool retrieval capabilities.
Downstream Applications. We also evaluate downstream applications with domain-specific toolsets: ALFWorld (Shridhar et al., 2021) (text embodied tasks with nine actions, e.g., move/take), WebShop (Yao et al., 2022a) (shopping with ‘search’ and ‘click’), GAIA (Mialon et al., 2024) (web search/browsing, VQA, code, file reading), and Humanity’s Last Exam (HLE) (Phan et al., 2025) (code, search, browsing, VQA). These tasks test long-horizon interaction in more realistic environments, requiring state tracking, error recovery, and coordination across heterogeneous tools; we equip agents with task-specific toolsets.
@@ 4.2. Baselines
Our baselines include: (1) Workflow-based Methods: ReAct (Yao et al., 2022b) alternates explicit reasoning with environment actions in a Reason-Act-Observe loop. CodeAct (Wang et al., 2024a) expresses actions as executable Python code that runs in an interpreter. Plan-and-Solve (Wang et al., 2023) first sketches a high-level plan and then executes it step by step. Reflexion (Shinn et al., 2023) enhances learning through verbal self-reflection after failed attempts. AgentLM (Zeng et al., 2024) uses instruction tuning to enhance general agent capabilities of LLMs. (2) Autonomous Tool Usage within Reasoning: WebThinker (Li et al., 2025b) interleaves thinking with web search and deep web exploration. HiRA (Jin et al., 2025c) introduces a hierarchical agent architecture where a meta planner decomposes tasks, a coordinator routes subtasks, and specialized executors solve them with dual-channel memory. OpenAI Deep Research (OpenAI, 2025) is an agentic system based on reasoning models.